# Big Tech influence over AI research revisited: memetic analysis of attribution of ideas to affiliation

### v1

@@ Abstract
Our findings suggest, that while Big Tech-affiliated papers are disproportionately more cited in some areas, the most cited papers are those affiliated with both Big Tech and Academia. Focusing on the most contagious memes, their attribution to specific affiliation groups (Big Tech, Academia, mixed affiliation) seems to be equally distributed between those three groups. This suggests that the notion of Big Tech domination over AI research is oversimplified in the discourse.
Ultimately, this more nuanced understanding of Big Tech’s and Academia’s influence could inform a more symbiotic alliance between these stakeholders which would better serve the dual goals of societal welfare and the scientific integrity of AI research.
@@ 1 Introduction
Artificial Intelligence (AI) research is often seen as dominated by Big Tech companies. These companies, with their immense computing resources and access to vast amounts of data, have undoubtedly influenced the development of the field. However, while this heavy influence has been beneficial in certain aspects, it has also raised concerns that the strong position of Big Tech could influence the direction and character of AI, resulting in significant losses for society and science. One of the issues raised is the reproducibility of Big Tech’s research (Ebell et al., 2021). Given their proprietary datasets and custom software, it is often difficult to independently replicate their studies. Conflict of interest is another issue that has been raised by researchers (Abdalla and Abdalla, 2021; Young et al., 2022; Hagendorff and Meding, 2021). One notable case that exposed these concerns was that of Dr. Timnit Gebru, formerly part of Google’s AI ethics team. Her dismissal highlighted the conflicts of interest that can arise when industry-aligned research agendas collide with ethical considerations. There is also evidence that the longer researchers remain in the industry, the more their high-impact work tends to become privatized, limiting the dissemination of knowledge that could otherwise benefit the academic community (Jurowetzki et al., 2021).
While there are many concerns about potential threats to research integrity due to this influence, it is equally important to recognize the positive impact of Big Tech companies. First and foremost, their computational resources and data, previously inaccessible on such a scale, provide unprecedented opportunities for the discovery and development of novel Machine Learning (ML) algorithms. In addition, Big Tech–Academia collaboration may benefit scientific endeavors for reasons other than access to resources. Evans (2010) argues that the industry’s relative lack of interest, in theory, could result in their academic partners being more likely to produce novel and theoretically unexpected experiments than academics working by themselves.
@@
Unfortunately, approaches to this topic mentioned above are limited. The methods used so far to quantify Big Tech’s dominance tend to oversimplify the concept of influence by perceiving it as the share of affiliations in papers regarding a specific topic. Methods are limited to tools such as topic modeling and keyword analysis. Current research also suffers from a lack of diversity in data sources, relying heavily on preprint servers such as arXiv or specific conferences to collect papers. In the previous studies mentioned above, the association of papers with affiliations was binary. That is, each paper was associated with either Big Tech (or the private sector in general) or Academia, depending on the first author or the proportion of affiliations in the paper. This limits conclusions that could be drawn from studies, as the Company–Academia collaboration in the area of AI is widespread.
All of the above limitations, combined with strong statements (e.g. researchers on this topic do not shy away from comparing Big Tech to Big Tobacco, Abdalla and Abdalla, 2021), could result in a limited, and possibly biased view of the matter, which could increase the tension between Big Tech and Academia, resulting in losses for science and society. Through our study, we aim to facilitate a more nuanced understanding of the dynamics between Big Tech and Academia and to highlight potential strengths and weaknesses within this relationship. We critically evaluate existing methodologies, advocate for more holistic measures of influence, and explore the implications of this dominance for the field of AI. We hope that this work will foster further discourse about the symbiosis between Academia and industry, and help to shape a more balanced AI research landscape.
In this study, along with a standard network analysis aimed at quantifying the influence of Big Tech and Academia on the AI papers citation network, we performed a memetic analysis. A meme is understood here as a piece of an idea that is transmitted through culture. In the context of research papers, this transmission occurs through citation. Using meme score, which is the quantification of the replicative power of a specific meme, we measure the spread of specific ideas in AI research. In addition, we quantify the probability of specific memes being replicated, conditioned on the affiliation of the authors of the paper containing the meme. This approach allows us to understand how the "spreading power" of specific ideas depends on affiliation groups (Big Tech, Academia, Companies, and mixed), which sheds light on which ideas Big Tech has more influence on spreading. Moreover, we investigated how papers with joint Big Tech–Academia affiliations differ from papers authored purely by authors affiliated with one category. In summary, we address the following research questions:
@@
Do papers affiliated with Companies and papers affiliated with both Big Tech and Academia differ in citation distributions from papers affiliated only with Big Tech or Academia?
@@ 2.1 Big tech influence over AI research
Several studies have attempted to quantify Big Tech’s influence in AI research, using a variety of methods and data sources. Abdalla and Abdalla (2021) examine the funding provided by Big Tech to academic AI researchers. The study identifies a recurring pattern where private companies increase their support for academic institutions when their public image declines, often due to media incidents such as the Cambridge Analytica scandal. The authors compare this funding pattern to the tactics employed by the tobacco industry. The study reveals that 59% of papers published in top journals that address the ethical and societal implications of AI include at least one author with financial ties to a Big Tech company. Ahmed and Wahed (2020) examine the changing dynamics of participation in major AI conferences following the rise of deep learning in 2012. The authors analyze 171,394 papers from 57 computer science conferences. They find an increase in the participation of large technology companies, particularly since 2012. In addition, the paper uses term frequency analysis of abstracts to uncover distinct research areas among different organizations.
@@
Zhang et al. (2019) employs altmetrics to assess the influence of AI publications, revealing an increased public interest in AI research findings since 2011. However, the literature suggests that altmetrics may not be suited to measuring societal impact Bornmann (2014).
@@ 2.1.1 Operationalization of the concept of Big Tech papers
Another approach is to look separately at papers that result from collaborations between Big Tech companies and other institutions. Articles describing such collaborations tend to have different characteristics in terms of format and tone compared to the research papers mentioned above, and focus on the positive aspects of such collaborations (Popkin, 2019). In recent work, (Färber and Tampakis, 2023) separatly analyzed papers co-authored by Academia and company-affiliated researchers. To the best of our knowledge, no analyses have specifically examined the influence and topics of interest reflected in papers resulting from collaborations between Big Tech and academia.
@@ 2.2 Memetic analysis
Several measures can characterize the spread of ideas in science (Wagner et al., 2011; Xu et al., 2018) or in business researcg (Wu et al., 2017). A straightforward method to identify ideas within a set of documents is to examine their frequency. Techniques like tf-idf (term frequency–inverse document frequency) allow the discovery of potential keywords. Topic modeling methods, such as Latent Dirichlet Allocation (LDA) or BERTopic (Grootendorst, 2022), help to understand the distribution of ideas consisting of multiple phrases across a document set. Another option is to focus on topical clusters, which can be built from bibliographic coupling and co-citation networks (Liu et al., 2017, 2019). One can use a citations cascade (Min et al., 2021) or chains of citations (della Briotta Parolo et al., 2020), which reveal a higher-order (i.e., reaching further than just the neighborhood of the node) temporal relationship between the papers.
@@ 3.1 Meme score
The sparking factor can be interpreted as the probability of the meme appearing without being present in any of the cited papers - how likely it is that the meme will be mentioned in the abstract, given that it does not appear in any of the cited papers’ abstracts. It is calculated by dividing the number of papers with the given meme that do not cite any papers with this meme by - all papers that do not cite any papers with the given meme.
@@
where is the number of papers with the meme , is the number of all memes in the dataset. is the number of papers that contain the meme and cite at least one paper with the meme and is the number of papers that cite at least one paper with the meme. is the number of papers containing the meme, but not citing a paper with the meme and is the number of papers that do not cite a paper with the meme.
From the three components of the meme score it is mainly the sticking factor that measures meme’s contagiousness. The other factors control if the phrase is not merely used as a part of a natural language or if a phrase is popular enough. Due to this, the sticking factor on its own could serve as a weaker measure of contagiousness.
@@ 3.2 Conditioned sticking factor
To measure how probable a meme is to spread under a given affiliation, we introduce a conditioned sticking factor, which is a modified version of the meme score introduced by Kuhn. A similar approach was proposed in the paper of Araújo and Fontainha (2018), where the authors attempted to condition the meme score on the main author’s gender. In contrast, in our approach, we condition only the sticking factor on the authors’ affiliations. The conditioned sticking factor is calculated as
@@
In contrast to the earlier approach of Araújo and Fontainha (2018), we have decided not to condition the sparking factor. The original purpose of the sparking factor is to control how often the word appears without any influence and limit the appearance of the most popular words that do not convey specific ideas. The sparking factor, conditioned analogously to the sticking factor, would correspond to the probability that the meme appears in a paper without being present in any cited papers with a given affiliation. However, the meme can appear in the other cited papers with different affiliations. In this way, the conditioned sparking factor would lose its original purpose of excluding phrases not conveying ideas without providing us with any specific additional information. For this reason, we decided to condition only the sticking factor. Finally, once , i.e., the meme is relatively common but rarely transmitted through the citation network, the meme score can be approximated by the sticking factor. Due to this, we focus primarily on the relationships between conditioned sticking factors for different groups. Figure 1 presents a toy example to explain how the sticking factor, sparking factor, meme score and conditioned sticking factor are calculated.
@@ 3.3 Papers dataset and processing
We gathered 557 681 articles from the S2ORC database, by searching for phrases proposed by Liu et al. (2021) (for details see A) in the titles and abstracts of the papers (all in English). We then performed exploratory data analysis. We found out that 138 878 records had no DOI identifiers, which we used to match the papers against the OpenAlex database. These papers were removed from our dataset. Next, we discovered that 146 525 records had no information about citations – such papers like this were useless for citation network analysis, so we decided to remove them from our dataset. Finally, we ended up with 166 455 records that had affiliations matched with OpenAlex.
OpenAlex provides information about the type of institution – it categorizes them into the following categories: company, education, government, facility, healthcare, nonprofit, and other (Table 2). By analyzing these categories, we decided to merge education, facility, and government into Academia. We also analyzed the company category and decided to label 14 companies with the largest number of publications as Big Tech.
The whole pipeline of memes extraction is shown in Fig. 10. All used abstracts were preprocessed first – we lowered all words and tokenized them. For the extraction of noun chunks, we used the spaCy library with the en_core_web_md model.
3.3.1 Papers’ affiliation
First, we look at the category of the company as it is specified in the OpenAlex Database. As Big Tech, we take a subset of companies that were among the top 60 companies by market capitalization in the years 2016-2019. The following companies fall into this category: Apple, Google, Microsoft, Facebook, Tencent, Oracle, Intel, IBM, Cisco Systems, TSMC, SAP, Qualcomm, Amazon, Siemens, Alibaba Group, Nvidia, and Samsung. Papers affiliated with Big Tech are 38% of the papers affiliated with the company category, but they account for 57% of their citations.
@@ 4.1 Non-binary distinction between Big Tech and academy–affiliated papers
We provide substantial arguments, advocating for the less coarse distinction between Academia and Big Tech affiliations than the bisection often used in the literature. The most natural way to follow this proposition would be to use a continuous space by simply calculating for each paper the proportion of Big Tech affiliations to the total number of affiliations. However, as shown in Fig. 2 these values are unevenly distributed – the vast majority are simply academic papers (almost 98% of all papers in the examined dataset), around 0.5% belong only to the Big Tech class, and the remaining part is characterized by a mixed affiliation. The last group also suffers from obvious finite-size effects, i.e., due to a limited number of authors, fractions such as , and tend to occur more frequently than other values (see Fig. 2). To avoid the above obstacles and to be consistent with the observations presented in Fig. 2, we decided to use a ternary affiliation scheme, i.e., Academia–mixed Big Tech–Big Tech.
@@
Let us first explore the implications of the ternary division using simple network metrics: incoming degree and PageRank (Brin and Page, 1998). Both these measures can be used to quantify the importance of a node (paper). In the case of the in-degree, we simply tally the number of citations a paper has received, while PageRank provides more nuanced information. Figure 3 shows box plots of in-degree (panel a) and PageRank (panel d), respectively, grouped according to the introduced ternary affiliation classification. In both cases, the Kruskal-Wallis test (Kruskal and Wallis, 1952) finds significant differences among all groups (p-value ), while post-hoc pairwise comparisons using Dunn’s test (Dunn, 1964) with Holm-Bonferroni corrections (Holm, 1979) indicate that the described distinctions are maintained for all pairs of groups. To prove that the division into three categories does not lead to a loss of additional information available in the continuous approach, we divide the dataset into six categories (Fig. 3c and Fig. 3f): the first and the last are the same as in the previous division, while the remaining four describe cumulative quartiles without exclusive Big Tech affiliations (note that the category labeled as “>0%” in Fig. 3c,f is equivalent to “mixed Big Tech” in Fig. 3a,d). Indeed, also at this level, we obtain similar results for the Kruskal-Wallis test (significant differences with p-value ) as well as for pairwise comparisons, the only exception being the penultimate category (“>75%”), which is indistinguishable from the pure Big Tech affiliations in the case of PageRank.
@@
To examine the reasons for the clear differences between the elements of our ternary classification, we focus on the in-degree and PageRank probability distributions shown in Fig. 3b and Fig. 3e, respectively. Clearly, mixed and Big Tech groups differ from Academia by being characterized by higher values of probability density for larger values of in-degree or PageRank. On the other hand, the difference between mixed and Big Tech is not as pronounced. The likely main reason for the observed differences is revealed when we decide to focus on the number of zero-degree nodes – the proportion of such nodes in Academia, mixed and Big Tech is 0.648, 0.538, and 0.613, respectively. Although the ratio of zero-degree nodes in the mixed category is visibly lower than in the other two categories, we confirm this observation by performing pairwise two-sample proportion tests with Holm-Bonferroni adjustment, which give p-values of for Academia–mixed Big Tech, for Academia – Big Tech, and for mixed Big Tech– Big Tech. An equivalent of Fig. 3 for the network with removed zero-degree (isolated) nodes is shown in D.
@@
Lastly, the decision to introduce the Big Tech category instead of looking at the OpenAlex company category was supported by the intuition that Big Techs differ significantly from other company actors and that extrapolating conclusions from all private companies to Big Tech would be unjustified. To test whether Big Tech are indeed different from the rest of the private companies, we compare the PageRanks of Big Tech–affiliated and company-affiliated papers. We perform the Kruskal-Wallis test between the PageRanks of Big Tech–affiliated papers and papers affiliated with companies that are not Big Tech. The test results confirm that the distribution of PageRanks of Big Tech-affiliated papers is indeed different from that of other companies (p-value < 0.05). Nevertheless, analyzing the ternary division from the company’s point of view (instead of Big Tech) does not bring any additional insights, as can be seen in Fig. 4 (see also D for results obtained with zero-degree nodes removed).

### v2

@@ Abstract
Our findings suggest that while Big Tech-affiliated papers are disproportionately more cited in some areas, the most cited papers are those affiliated with both Big Tech and Academia. Focusing on the most contagious memes, their attribution to specific affiliation groups (Big Tech, Academia, mixed affiliation) seems equally distributed between those three groups. This suggests that the notion of Big Tech domination over AI research is oversimplified in the discourse.
@@ 1 Introduction
Artificial Intelligence (AI) research is often seen as dominated by Big Tech companies. These companies, with their immense computing resources and access to vast amounts of data, have influenced the development of the field. However, while this influence has been beneficial in certain aspects, it has also raised concerns that the strong position of Big Tech could influence the direction and character of AI, resulting in significant losses for society and science. One of the issues raised is the reproducibility of Big Tech’s research (Ebell et al., 2021). Given their proprietary datasets and custom software, independently replicating their studies is often difficult. Conflict of interest is another issue that has been raised by researchers (Abdalla and Abdalla, 2021; Young et al., 2022; Hagendorff and Meding, 2021). One notable case that exposed these concerns was that of Dr. Timnit Gebru, formerly part of Google’s AI ethics team. Her dismissal highlighted the conflicts of interest arising when industry-aligned research agendas collide with ethical considerations. There is also evidence that the longer researchers remain in the industry, the more their high-impact work becomes privatized, limiting the dissemination of knowledge that could otherwise benefit the academic community (Jurowetzki et al., 2021).
While there are many concerns about potential threats to research integrity due to this influence, it is equally important to recognize the positive impact of Big Techs . First, their computational resources and data, previously inaccessible on such a scale, provide unprecedented opportunities to discover and develop novel Machine Learning (ML) algorithms. In addition, Big Tech–Academia collaboration may benefit scientific endeavors for reasons other than access to resources. Evans (2010) argues that the industry’s relative lack of interest, in theory, could result in their academic partners being more likely to produce novel and theoretically unexpected experiments than academics working by themselves.
@@
Unfortunately, approaches to this topic mentioned above are limited. The methods used so far to quantify Big Tech’s dominance tend to oversimplify the concept of influence by perceiving it as the share of affiliations in papers regarding a specific topic. Methods are limited to tools such as topic modeling and keyword analysis. Current research also lacks diversity in data sources, relying heavily on preprint servers or specific conferences to collect papers. In the previous studies, the association of papers with affiliations was binary – each paper was associated with either Big Tech (or the private sector in general) or Academia, depending on the first author or the proportion of affiliations in the paper. This limits possible conclusions, as the Company–Academia collaboration in the area of AI is widespread.
All of the above limitations, combined with strong statements (e.g. researchers on this topic do not shy away from comparing Big Tech to Big Tobacco, Abdalla and Abdalla, 2021), could result in a selective view of the matter, leading to the tension between Big Tech and Academia, and losses for science and society. Through our study, we aim to facilitate a more nuanced understanding of Big Tech and Academia dynamics, highlighting strengths and weaknesses within this relationship. We critically evaluate existing methodologies, advocate for more holistic measures of influence, and explore the implications of this dominance for the field of AI to foster further discourse about the symbiosis between Academia and industry.
In this study, along with a network analysis aimed at quantifying the influence of Big Tech and Academia on the AI papers citation network, we performed memetic analysis. A meme is understood here as a piece of an idea that is transmitted through culture. In the context of research papers, this transmission occurs through citation. Using meme score, which quantifies the replicating power of a specific meme, we measure the spread of particular ideas in AI research. In addition, we quantify the probability of particular memes being replicated, conditioned on the affiliation of the authors of the paper containing the meme. This approach allows us to understand how the “spreading power” of specific ideas depends on affiliation groups (Big Tech, Academia, Companies, and mixed), which sheds light on which ideas Big Tech has more influence on spreading. Moreover, we investigated how papers with joint Big Tech–Academia affiliations differ from papers authored purely by authors affiliated with one category. In summary, we address the following research questions:
@@
Do papers affiliated with both Big Tech and Academia differ in citation distributions from papers affiliated only with Big Tech or Academia? Do Big Tech affiliations differ from Company ones (i.e., industrial but not attributed to Big Tech)?
@@ 2.1 Big tech influence over AI research
Several studies have attempted to quantify Big Tech’s influence in AI research using a variety of methods and data sources. Abdalla and Abdalla (2021) examine the funding provided by Big Tech to academic AI researchers. The study identifies a recurring pattern where private companies increase their support for academic institutions when their public image declines, often due to media incidents such as the Cambridge Analytica scandal. The authors compare this funding pattern to the tactics employed by the tobacco industry. The study reveals that 59% of papers published in top journals that address AI’s ethical and societal implications include at least one author with financial ties to a Big Tech company.
Ahmed and Wahed (2020) examine the dynamics of participation in major AI conferences following the rise of deep learning in 2012. The authors analyze 171,394 papers from 57 computer science conferences. They find an increase in the participation of large technology companies, particularly since 2012. In addition, the paper uses term frequency analysis of abstracts to uncover distinct research areas among different organizations.
@@
Zhang et al. (2019) employs altmetrics to assess the influence of AI publications, revealing an increased public interest in AI research findings since 2011. However, the literature suggests that altmetrics may not be suited to measuring societal impact (Bornmann, 2014).
@@ 2.1.1 Operationalization of the concept of Big Tech papers
Another approach is to look separately at papers that result from collaborations between Big Tech companies and other institutions. Articles describing such collaborations tend to have different characteristics in terms of format and tone compared to the research papers mentioned above, and focus on the positive aspects of such collaborations (Popkin, 2019). In recent work, (Färber and Tampakis, 2023) separately analyzed papers co-authored by Academia and company-affiliated researchers. To the best of our knowledge, no analyses have specifically examined the influence and topics of interest reflected in papers resulting from collaborations between Big Tech and academia.
@@ 2.2 Memetic analysis
Several measures can characterize the spread of ideas in science (Wagner et al., 2011; Xu et al., 2018) or in business research (Wu et al., 2017). A straightforward method to identify ideas within a set of documents is to examine their frequency. Techniques like tf-idf (term frequency–inverse document frequency) allow the discovery of potential keywords. Topic modeling methods, such as Latent Dirichlet Allocation (LDA) or BERTopic (Grootendorst, 2022), help to understand the distribution of ideas consisting of multiple phrases across a document set. Another option is to focus on topical clusters, which can be built from bibliographic coupling and co-citation networks (Liu et al., 2017, 2019). One can use a citations cascade (Min et al., 2021) or chains of citations (della Briotta Parolo et al., 2020), which reveal a higher-order (i.e., reaching further than just the neighborhood of the node) temporal relationship between the papers.
@@ 3.1 Meme score
The sparking factor can be interpreted as the probability of the meme appearing without being present in any of the cited papers - how likely it is that the meme will be mentioned in the abstract, given that it does not appear in any of the cited papers’ abstracts. It is calculated by dividing the number of papers with the given meme that do not cite any papers with this meme by all papers that do not cite any papers with the given meme.
@@
where is the number of papers with the meme , is the number of all memes in the dataset. is the number of papers that contain the meme and cite at least one paper with the meme and is the number of papers that cite at least one paper with the meme. On the other hand is the number of papers containing the meme but not citing a paper with the meme, and is the number of papers that do not cite a paper with the given meme.
From the three components of the meme score, it is mainly the sticking factor that measures the meme’s contagiousness. The other factors control if the phrase is not merely used as a part of a natural language or if a phrase is popular enough. Due to this, the sticking factor on its own could serve as a weaker measure of contagiousness. Meme score can be simply interpreted as a ratio of the probability of replicating a phrase to the probability of its appearance without earlier presence in the cited papers. To select meaningful memes, we must choose a threshold of the meme score defining which phrases are considered memes and which are not. This is done by selecting the point of maximum curvature from the value of the meme score as a function of its position in the phrase ranking ordered by the meme score (see Sec. 4.2 for practical implementation of the referred method). Such an approach is similar to the so-called “elbow method” (Ketchen and Shook, 1996) – a commonly used visual technique in assessing the optimal number of clusters in the k-means method.
@@ 3.2 Conditioned sticking factor
To measure how probable a meme is to spread under a given affiliation, we introduce a conditioned sticking factor – a modified version Kuhn’s meme score. A similar approach was proposed in Araújo and Fontainha (2018), where the authors conditioned the meme score on the main author’s gender. In contrast, we condition only the sticking factor on the authors’ affiliations. The conditioned sticking factor is calculated as
@@
In contrast to the earlier approach (Araújo and Fontainha, 2018), we have decided not to condition the sparking factor. The original purpose of the sparking factor is to control how often the word appears without any influence and limit the appearance of the most popular words that do not convey specific ideas. The sparking factor, conditioned analogously to the sticking factor, would correspond to the probability that the meme appears in a paper without being present in any cited papers with a given affiliation. However, the meme can appear in the other cited papers with different affiliations. In this way, the conditioned sparking factor would lose its original purpose of excluding phrases not conveying ideas without providing us with any specific additional information. For this reason, we decided to condition only the sticking factor. Finally, once , i.e., the meme is relatively common but rarely transmitted through the citation network, the meme score can be approximated by the sticking factor. Due to this, we focus primarily on the relationships between conditioned sticking factors for different groups. Figure 1 presents a toy example to explain how the sticking factor, sparking factor, meme score, and conditioned sticking factor are calculated.
@@ 3.3 Papers dataset and processing
As a main data source for this study, we used the S2ORC database Lo et al. (2020), which is a large corpus of 81.1 million English-language academic papers spanning many disciplines. The collection comes from various sources, such as in-proceedings of scientific conferences, well-established journals, and digital archives (e.g. arXiv). We gathered 557 681 articles from this source by searching for phrases presented in A, which occurred in the titles and abstracts of the papers. The aforementioned set of keywords is based on the list proposed by Liu et al. (2021), who conducted the study, attempting to find a representative set of keywords for the AI domain while maintaining a reasonable balance between recall and precision (we consider possible data limitations in the Discussion section).
Afterwards, we performed exploratory data analysis. We found out that 138 878 records had no DOI identifiers, which we used to match the papers against the OpenAlex database. As affiliation data provided by those links is crucial for our study, we decided to remove such papers from our dataset. Next, we discovered that 146 525 of the remaining records had no information about citations. Papers like this were useless for citation network analysis, so we decided to remove them from our dataset. Finally, we ended up with 166 455 records that had affiliations matched with OpenAlex.
The whole pipeline of memes extraction is shown in Fig. 10. All used abstracts were preprocessed first – we lowercased all words and tokenized them. For the extraction of noun chunks, we used the spaCy library with the en_core_web_md model.
Papers’ affiliation
First, we look at the category of the company as it is specified in the OpenAlex Database. OpenAlex provides information about the type of institution – it categorizes them into the following categories: company, education, government, facility, healthcare, nonprofit, and other (Table 2). By analyzing these categories, we decided to merge education, facility, and government into Academia. We also analyzed the company category separately.
As Big Tech, we take a subset of technological companies that were among the top 60 companies by market capitalization in the years 2016-2019 (PwC, 2019, 2018, 2017, 2016). The following companies fall into this category: Apple, Google, Microsoft, Facebook, Tencent, Oracle, Intel, IBM, Cisco Systems, TSMC, SAP, Qualcomm, Amazon, Siemens, Alibaba Group, Nvidia, and Samsung. We aimed to construct a list that encompasses companies previously labelled as Big Techs in media, such as the Big Five, FAANG or MAGA (Economist, 2018), while also being supported by a quantitative measure like market capitalization. To ensure the list’s robustness and account for short-term market changes, we use a union of companies that appeared in the top 60 companies during the four years with the highest number of papers in our dataset. We opt to limit ourselves to the top 60 companies to include companies considered most significant to the field while excluding the lower end of the top 100 companies, where most technological companies appeared only once during the four years we consider. Papers affiliated with Big Tech are 38% of the papers affiliated with the company category, but they account for 57% of their citations.
@@ 4.1 Non-binary distinction between Big Tech and academy–affiliated papers
We provide substantial arguments, advocating for the less coarse distinction between Academia and Big Tech affiliations than the bisection often used in the literature. The most natural way to follow this proposition would be to use a continuous space by simply calculating for each paper the proportion of Big Tech affiliations to the total number of affiliations. However, as shown in Fig. 2, these values are unevenly distributed – the vast majority are simply academic papers (almost 98% of all papers in the examined dataset), around 0.5% belong only to the Big Tech class, and the remaining part is characterized by a mixed affiliation. The last group also suffers from obvious finite-size effects, i.e., due to a limited number of authors, fractions such as , and tend to occur more frequently than other values (see Fig. 2). To avoid the above obstacles and to be consistent with the observations presented in Fig. 2, we decided to use a ternary affiliation scheme, i.e., Academia–mixed Big Tech–Big Tech.
@@
Let us first explore the implications of the ternary division using simple network metrics: incoming degree and PageRank (Brin and Page, 1998). Both these measures can be used to quantify the importance of a node (paper). In the case of the in-degree, we simply tally the number of citations a paper has received, while PageRank provides more nuanced information. Figure 3 shows in-degree (panel a) and PageRank (panel c) distributions, grouped according to the introduced ternary affiliation classification (denoted as “Academia”, “Mixed” and “Big Tech”). In both cases, the Kruskal-Wallis test finds significant differences among all groups (p-value ), while post-hoc pairwise comparisons using Dunn’s test with Holm-Bonferroni corrections indicate that the described distinctions are maintained for all pairs of groups (if two or more categories are statistically indistinguishable, they are connected with a solid line on the far right part of the panel). To prove that the division into three categories does not lead to a loss of information available in the continuous approach, we divide the dataset into six categories: the first two and the last one are the same as in the previous division, while the remaining three describe cumulative quartiles without exclusive Big Tech affiliations (“”, “” and “”, i.e., “” means that each node in this category has over 1/4 share of Big Tech affiliations but cannot be exclusively Big Tech). We obtain similar results for the Kruskal-Wallis test (significant differences with p-value ) as well as for pairwise comparisons, the only exception being the penultimate category (“”), which is indistinguishable from the pure Big Tech affiliations in the case of PageRank. All the intermediate categories (“Mixed”, “”, “” and “”) are indistinguishable from each other, proving that further divisions in the “Mixed” category lack additional insights.
@@
To examine the reasons for the clear differences between the elements of our ternary classification, we focus on the in-degree and PageRank probability distributions shown in Fig. 3b and Fig. 3d, respectively. Clearly, mixed and Big Tech groups differ from Academia by higher probability density values for larger values of in-degree or PageRank. On the other hand, the difference between mixed and Big Tech is not as pronounced. The likely reason for the observed differences is connected to the number of zero-degree nodes – the proportion of such nodes in Academia, mixed and Big Tech is 0.648, 0.538, and 0.613, respectively. Although the ratio of zero-degree nodes in the mixed category is visibly lower than in the other two categories, we confirm this observation by performing pairwise two-sample proportion tests with Holm-Bonferroni adjustment, which give p-values of for Academia–mixed Big Tech, for Academia – Big Tech, and for mixed Big Tech– Big Tech. Thus, the network with removed zero-degree nodes brings different statistical results marked in Fig. 3a and Fig. 3c by the dotted line on the far right – in this case, we recover the binary division as only “Academia” category is distinguishable from any other.
@@
Lastly, the decision to introduce the Big Tech category instead of using OpenAlex company category was supported by the intuition that Big Techs differ significantly from other company actors and that extrapolating conclusions from all private companies to Big Tech would be unjustified. To test whether Big Techs are different from the rest of the private companies, we compare the PageRanks of Big Tech–affiliated and company-affiliated papers, performing the Kruskal-Wallis test between the PageRanks of Big Tech–affiliated papers and papers affiliated with companies that are not Big Tech. The results confirm that the distribution of PageRanks of Big Tech-affiliated papers is indeed different from that of other companies (p-value ¡ 0.05). Nevertheless, analyzing the ternary division from the company’s point of view (instead of Big Tech) does not bring any additional insights, as can be seen in Fig. 4.